OpenStudy (anonymous):

what is the molar mass of 3Na2CO3

OpenStudy (joannablackwelder):

This is kind of a funky question because the 3 out in front of the compound means 3 moles, but molar mass is, by definition, per 1 mole.

OpenStudy (joannablackwelder):

So, I think you can just find the molar mass of Na2CO3.

OpenStudy (joannablackwelder):

You can do that by adding up the atomic masses of each of the elements, multiplied by how many of that element you have in 1 compound.

OpenStudy (anonymous):

ah sorry. here is the original question: If 852.04 g Al(NO3)3 reacted with the 741.93 g Na2CO3, which would be the limiting reagent? (2 points)

OpenStudy (anonymous):

2 Al(NO3)3 + 3 Na2CO3 = Al2(CO3)3(s) + 6 NaNO3

OpenStudy (joannablackwelder):

Ah, I see. I would use conversions to find how much of 1 of the products you can make with each amount of reactant you have.

OpenStudy (joannablackwelder):

The reactant that makes less of the same product will run out first, or be the limitingreactant.

OpenStudy (anonymous):

yep. i was finding the masses of each

OpenStudy (anonymous):

and ratios

OpenStudy (joannablackwelder):

Sounds good. :)

OpenStudy (anonymous):

but I'm not sure of the mass of 3Na2CO3

OpenStudy (anonymous):

the answer is 341 but i can't get that

OpenStudy (joannablackwelder):

I don't get that either...

OpenStudy (joannablackwelder):

I get the molar mass to be 105.98g/mol

OpenStudy (anonymous):

i get 236

OpenStudy (anonymous):

so the mole ratios are 2, 3, 1, and 6

OpenStudy (anonymous):

and the molar mass of 2Al(NO3)3 is 240.034

OpenStudy (joannablackwelder):

To find the molar mass, you just use the subscript numbers. 2 Na, 1 C, and 3 O.

OpenStudy (joannablackwelder):

For. Na2CO3.

OpenStudy (joannablackwelder):

You can multiply that up to find the mass of 3 moles of the compound, but that is not the molar mass.

OpenStudy (anonymous):

ah okay

OpenStudy (joannablackwelder):

:)

OpenStudy (anonymous):

okay i really need help. so i figured out that

OpenStudy (anonymous):

7.00 moles of Na2CO3 4.00 moles of Al(NO3)3 are needed

OpenStudy (anonymous):

Al(NO3)3 : k1 = n(Al(NO3)3) / 2 = 4,00 / 2 = 2,00 Na2CO3 : k2 = n(Na2CO3) / 3 = 7,00 / 3 = 2,33 OR 7 mol of Na2CO3 times 2 mole of Al(NO3) divided by 3 mol of NA2COS

OpenStudy (joannablackwelder):

All your calculations look great to me! So, what does it mean?